Sir,The final velocity of a particle falling freely under gravity is given by V^2-u^2=2gx where x is the distance covered. If V=18kmph g=1000cms^-2 x=120cm the u = ms^-1

Grade:11

1 Answers

Vikas TU
7 years ago
Just put all the values simply to get the value of u.
Therefore,
V^2 – u^2 = 2*g*x
(18 * 1000 * 100/60*60)^2 = u^2 + 2*1000*120
u^2 = 250000 – 240000
     u^2 = 10,000
 u =  100 cm/s